How many different committees can be formed from 5 professors and 15 students if each committee is made up of 2 professors and 10 students?
Answer by stanbon(75887)

How many different committees can be formed from 5 professors and 15 students if each committee is made up of 2 professors and 10 students?
------------------------------------
5C2 * 15C10 = 10*3003 = 30030
======================================
Cheers,
Stan H.